solve using addition method:
3x^2-2y^2=-5
2x^2-y^2=-2

Answer by checkley77(12844)

3x^2-2y^2=-5
2x^2-y^2=-2 multiply this equation by -2 & add.
-4x^2+2y^2=4
3x^2-2y^2=-5
-------------------
x^2+0=-1
x^2+1=0
x=0+/-1i (this has complex roots)